1. Valles de Benasque y Eriste

FORCAU OR ANGEL ORÚS REFUGE 2 hours

Mountain refuge with forest guard in the Grist Valley (Eriste) where you can spend the night and have some meals. This refuge is the base for the ascent to Posets and some other peaks in the valley.

Once in Grist (Eriste), take the road that goes up the slope from the town called Aragonesas, next to the power station. The road joins a forest trail (usually in good condition) that goes into the Grist valley. After 6 km. from the town, the trail ends and you have to go up through the woods where the valley begins. On your right a concrete brigde crosses the river: \underline{do not take it}. Take the track on the left of the valley until you get to the Cabana Plleta de Sallén (the hut at the Plleta de Sallén) after a 1 hour walk. From here the track go up the left slope with a steep incline till you get to the Plleta and Forcau refuge (or Angel Orús refuge) (2,045 m. high).